# Watch and Learn: Learning to Use Computers from Online Videos

### v2

@@ Abstract
1 Introduction
Computer-using agents (CUAs) (Zheng et al., 2024a; Kil et al., 2024; Qin et al., 2025; Gou et al., 2025; OpenAI, 2025b; Yang et al., 2025b; Wang et al., 2025; Xie et al., 2025; Agashe et al., 2025a, b; Gonzalez-Pumariega et al., 2025) are emerging as a powerful paradigm for automating software and web interactions. From everyday productivity tools to enterprise workflows, CUAs must both plan multi-step task sequences that incorporate domain knowledge and ground these plans into concrete UI actions across diverse, constantly changing interfaces. Despite rapid progress, their scalability remains constrained by the difficulty of obtaining large-scale, high-quality demonstrations of human-computer interaction (Wang et al., 2025).
@@ 1 Introduction
We evaluate W&L on two challenging benchmarks, OSWorld (Xie et al., 2024) and WindowsAgentArena (Bonatti et al., 2024), which require complex reasoning and precise UI grounding. On OSWorld, trajectories extracted from web videos yield consistent gains: in-context use improves general-purpose models and agentic frameworks by up to 3 percentage points, while supervised fine-tuning provides larger boosts for open-weight models (up to 11 points). On WindowsAgentArena, fine-tuning with our trajectories achieves leading performance among 7B-scale models under the 15-step evaluation limit, demonstrating that large-scale web demonstrations can translate effectively into high-quality training data. These benefits are obtained without any manual annotation, highlighting the potential of web-scale human workflows as a practical and scalable foundation for advancing CUAs toward real-world deployment. In summary, this work makes the following contributions:
@@ 4.1 Experimental Setup
Our experiments evaluate whether IDM-labeled trajectories improve interactive computer-using agents in two settings: (1) in-context learning (ICL) for general-purpose multimodal models and agentic frameworks, and (2) supervised fine-tuning (SFT) for open-source models. We also compare the resulting trajectories against TongUI (Zhang et al., 2025), an existing automatic labeling pipeline.
@@ Acknowledgments
We thank the Google Cloud AI Research team for their valuable feedback, with special thanks to Yale Song for his constructive input.
@@ D.1 Vision Encoder
Model Selection. We employ SigLIP-2 Base (patch size 16) with the NaFlex variant as our vision encoder [Tschannen et al., 2025]. This architecture was specifically chosen for its superior performance on OCR and document-understanding benchmarks, which closely aligns with the requirements of computer interface understanding where precise text recognition and spatial localization are critical.
@@
Rationale for NaFlex. Compared to the fixed-resolution SigLIP-2 variant at $\times512$ (which would downsample our $\times1000$ input by 48%), NaFlex provides: (1) Higher effective resolution ($\sim$$\times704$) for better preservation of fine-grained UI details, (2) Superior performance on text-heavy and structured visual content [Tschannen et al., 2025], and (3) Minimal aspect ratio distortion for square inputs. While fixed-resolution variants benefit from additional distillation stages, preliminary experiments showed NaFlex’s advantages in OCR-like capabilities outweigh this trade-off for UI understanding.
@@ D.3.3 Language Head
Architecture. For text-entry actions (type), we employ a lightweight autoregressive decoder based on GPT-2 [Radford et al., 2019].
@@ F.2 Which application domains benefit most from our data?
The following rules are important: If previous actions did not achieve the expected result, do not repeat them. Adjust your action or coordinates based on the new screenshot. Do not predict multiple clicks at once. Base each action only on what is currently visible. You cannot complete the task by outputting text. You must interact through mouse and keyboard. Return Fail when the task cannot be done. Observation. Provide a detailed observation of the current computer state based on the full screenshot in an “Observation:” section. Include any potentially relevant information such as pop-ups, notifications, or error messages. You must output the observation before the thought.
Combined ICL Prompt for o3 (continued) Thought. Think step by step and describe the full reasoning process behind your next action. Your thought should include: a progress assessment and reflection on previous attempts, identification of errors or mismatches, enumeration of possible next actions based on visible state, analysis and selection of the best next action. You must output the thought before the action code. Action Format. Use pyautogui to perform actions, but do not use pyautogui.locateCenterOnScreen or pyautogui.screenshot(). Return exactly one line of Python code, preceded by a clear instruction comment. Return the code inside a code block:
@@
Special tokens. WAIT: when a pause is needed, DONE: when the task is completed, FAIL: when the task cannot be completed. You have a maximum of 100 steps and the current step is {current_step} out of {max_steps}. My computer’s password is ’{CLIENT_PASSWORD}’. Use it when necessary for sudo operations. First, briefly reflect on the current screenshot and prior actions, then return only the code or special token.

### v3

@@ Abstract
Project page: https://chanh.ee/wandl/.
1 Introduction
Computer-using agents (CUAs) (Zheng et al., 2024a; Kil et al., 2024; Qin et al., 2025; Gou et al., 2025; OpenAI, 2025b; Yang et al., 2025b; Wang et al., 2025; Xie et al., 2025; Agashe et al., 2025a, b; Gonzalez-Pumariega et al., 2025) are emerging as a powerful paradigm for automating software and web interactions. From everyday productivity tools to enterprise workflows, CUAs must both plan multi-step task sequences that incorporate domain knowledge and ground these plans into concrete UI actions across diverse, constantly changing interfaces. Despite rapid progress, their scalability remains constrained by the difficulty of obtaining large-scale, high-quality demonstrations of human-computer interaction Wang et al. (2025).
@@ 1 Introduction
We evaluate W&L on two challenging benchmarks, WindowsAgentArena (Bonatti et al., 2024) and OSWorld (Xie et al., 2024), which require complex reasoning and precise UI grounding. On OSWorld, trajectories extracted from web videos yield consistent gains: in-context use improves general-purpose models and agentic frameworks by up to 3 percentage points, while supervised fine-tuning provides larger boosts for open-weight models (up to 11 points). On WindowsAgentArena, fine-tuning with our trajectories achieves leading performance among 7B-scale models under the 15-step evaluation limit, demonstrating that large-scale web demonstrations can translate effectively into high-quality training data. These benefits are obtained without any manual annotation, highlighting the potential of web-scale human workflows as a practical and scalable foundation for advancing CUAs toward real-world deployment. In summary, this work makes the following contributions:
@@ 4.1 Experimental Setup
Our experiments evaluate whether IDM-labeled trajectories improve interactive computer-using agents in two settings: (1) in-context learning (ICL) for general-purpose multimodal models and agentic frameworks, and (2) supervised fine-tuning (SFT) for open-source models. We also compare the resulting trajectories against TongUI Zhang et al. (2025), an existing automatic labeling pipeline.
@@ Acknowledgments
We thank the Google Cloud AI Research team for their valuable feedback, with special thanks to Yale Song for his constructive input. This work was supported in part by the National Science Foundation under NSF CAREER Award #2443149 and by an Alfred P. Sloan Foundation Fellowship.
@@ C.5 Models
C.6 Action Space and Primitive Composition
The inverse dynamics model (IDM) predicts atomic GUI interaction primitives from observed state transitions. The primitive action space consists of Click, Release, Move, Scroll, Type, and Wait. These primitives serve as a low-level interaction vocabulary and are not executed directly in the environment.
[…]
Primitive distinctions are introduced only when they induce observable changes in the visual state. When multiple interaction sequences are visually indistinguishable under the observation stream, they are resolved to the minimal primitive sequence consistent with the observed transition. This constraint avoids introducing latent action distinctions that cannot be inferred from visual evidence.
To ensure compatibility with downstream execution, the IDM primitive space is aligned with a subset of the agent and environment action spaces through the deterministic mapping described above. This alignment constrains the prediction space to executable interactions and mitigates degenerate solutions during training by preventing the model from collapsing onto non-executable or redundant action representations.
@@ D.1 Vision Encoder
Model Selection. We employ SigLIP-2 Base (patch size 16) with the NaFlex variant as our vision encoder Tschannen et al. [2025]. This architecture was specifically chosen for its superior performance on OCR and document-understanding benchmarks, which closely aligns with the requirements of computer interface understanding where precise text recognition and spatial localization are critical.
@@
Rationale for NaFlex. Compared to the fixed-resolution SigLIP-2 variant at $\times512$ (which would downsample our $\times1000$ input by 48%), NaFlex provides: (1) Higher effective resolution ($\sim$$\times704$) for better preservation of fine-grained UI details, (2) Superior performance on text-heavy and structured visual content Tschannen et al. [2025], and (3) Minimal aspect ratio distortion for square inputs. While fixed-resolution variants benefit from additional distillation stages, preliminary experiments showed NaFlex’s advantages in OCR-like capabilities outweigh this trade-off for UI understanding.
@@ D.3.3 Language Head
Architecture. For text-entry actions (type), we employ a lightweight autoregressive decoder based on GPT-2 Radford et al. [2019].
@@ F.2 Which application domains benefit most from our data?
The following rules are important: If previous actions did not achieve the expected result, do not repeat them. Adjust your action or coordinates based on the new screenshot. Do not predict multiple clicks at once. Base each action only on what is currently visible. You cannot complete the task by outputting text. You must interact through mouse and keyboard. Return Fail when the task cannot be done. Observation. Provide a detailed observation of the current computer state based on the full screenshot in an “Observation:” section. Include any potentially relevant information such as pop-ups, notifications, or error messages. You must output the observation before the thought. Thought. Think step by step and describe the full reasoning process behind your next action. Your thought should include: a progress assessment and reflection on previous attempts, identification of errors or mismatches, enumeration of possible next actions based on visible state, analysis and selection of the best next action. You must output the thought before the action code. Action Format. Use pyautogui to perform actions, but do not use pyautogui.locateCenterOnScreen or pyautogui.screenshot(). Return exactly one line of Python code, preceded by a clear instruction comment. Return the code inside a code block:
@@
Combined ICL Prompt for o3 (continued) Special tokens. WAIT: when a pause is needed, DONE: when the task is completed, FAIL: when the task cannot be completed. You have a maximum of 100 steps and the current step is {current_step} out of {max_steps}. My computer’s password is ’{CLIENT_PASSWORD}’. Use it when necessary for sudo operations. First, briefly reflect on the current screenshot and prior actions, then return only the code or special token.